# Coarse-graining nonequilibrium diffusions with Markov chains

### v1

@@ 5.2.1. An application to real-world trajectories: group-polarisation in schooling fish
Finally, we apply this approach to a real-world case of a stationary diffusion, the polarisation trajectories of schooling fish, and determine that such a process is not in a NESS. We follow the analysis of Jhawar et al [Jhawar2020schoolingfish], who analyse populations of $N=15,30$ and 60 fish moving in a tank. In particular, the positions, $\mathbf{x}_{i}(t)$, and velocities, $\mathbf{v}_{i}(t)$, of the fish can be extracted from video recordings. The group polarisation is an order parameter describing their collective alignment,
@@
Next, we investigate if the stationary dynamics of the schooling fish are best described as a NESS or an ESS. For each of the population sizes, we infer the EPR of the original trajectory as well as the EPR of 100 surrogate models, obtained using the shuffling procedure. Fig. 18 shows that, across population size, the EPR is not significantly higher for the original trajectories when compared to the surrogate model, from which we conclude that the process is not in a NESS. This result supports a well-known phenomenon in collective dynamics and active matter. Whilst collective dynamics are inherently active at a local level [toner1998flocks], as they dissipate energy in order for agents to adjust their states dynamically, their collective behaviour can be described by equilibrium steady-state dynamics respecting time-reversal symmetry [Fodor2016howfar] and maximum-entropy distributions [Bialek2012stat, Mora2016local].
@@ 6. Discussion
Nonequilibrium steady-states are ubiquitous phenomena in physical and biological systems [Gnesotto2018brokendetailedbalance, nartallokalu2025review, fang2019nonequilibrium]. Approximating the continuous-state dynamics of these diffusions with discrete-state models is a simple and effective way to analyse their dynamics, and has found applications to physical and biological data at multiple scales [battle2016brokendetailedbalance, Li2019quantifying, lynn2021detailedbalance, nartallokaluarachchi2024decomposing]. Nevertheless, the consistency of the properties of the NESS between the continuous and discrete spaces was not well understood. Moreover, these approaches suffer the complex and obscurant effects of coarse-graining which can hide probability flux leading to underestimates of the entropy production rate [Esposito2012coarsegraining] as well as induce non-Markovian effects that lead to artefactual scaling relationships [Schwarz2024Memory].
In this paper, we presented a new approach to analytically derive a discrete-state Markov process from stationary planar diffusion, using finite-volume methods. Moreover, we showed that the EPR of the approximation converges in the infinite-state limit, which we illustrated with both solvable and unsolvable examples of stationary diffusions. In the case of unsolvable processes, this approximation gives an accurate way to investigate the behaviour of the EPR under changes to the parameters, which we illustrated with the stochastic van der Pol and Kuramoto models. Finally, we considered the problem of statistical inference from trajectory data. Using sampled paths, our results show that discrete-state models significantly underestimate the EPR of a NESS, even for long trajectories and fine grids. Nevertheless, discrete-state models can be used to define a hypothesis test that can be used to determine if a trajectory is from a NESS or not, which we, again, illustrated with sampled trajectories. Finally, we considered an illustrative application to real-world trajectories by analysing the group-polarisation trajectories from schooling fish [Jhawar2020schoolingfish].
@@
Python code, Jupyter notebooks, and trajectory data will be made available at https://github.com/rnartallo/coarse_grained_diffusion. The full dataset for the schooling fish has been made available by Jhawar et al at https://zenodo.org/records/3632470.
@@ Acknowledgements
The authors would like to thank Alexander Strang, Tassilo Schwarz and Jean-Charles Delvenne for important conversations about coarse-graining and stochastic processes. Additionally, they would like to thank Endre Süli, José Antonio Carrillo and Maria Bruna for their advice on finite-volume discretisations. Finally, they would like to thank Jhawar et al for making their data on schooling fish publicly available.
@@ Appendix A The Scharfetter-Gummel discretisation
To derive the Scharfetter-Gummel discretisation in Eq. (25), we follow the analysis of Ref. [Frensley2004scharfettergummel], extending to the 2D case with non-constant drift. We begin with the FVA of the FP equation,
@@
which is Eq. (23) of the main text. Next, we assume that the flux $J^{x}_{i+\frac{1}{2}}$, drift $f^{x}_{i+\frac{1}{2}}$ and diffusion $D^{x}_{i+\frac{1}{2}}$ are constant over the interval $x_{i},x_{i+1}]$. As a result, $p^{j}(x):=p(x,y_{j})$ satisfies the first-order ODE with constant coefficients,
(71) $J^{x}_{i+\frac{1}{2}}$ $=f^{x}_{i+\frac{1}{2},j}p^{j}(x)-D^{x}_{i+\frac{1}{2},j}\frac{dp^{j}}{dx}(x),$
@@
We then integrate this over the range $x_{i},x_{i+1}]$ where $\Delta x=x_{i+1}-x_{i}$. This yields,
@@
Reintroducing the subscripts, we obtain,
@@
Whilst not detailed here, we found that other structure-preserving schemes such as the Chang-Cooper discretisation [Pareschi2018structure] yielded valid master equations for these diffusions, whilst the centred-difference scheme only did so under step-size conditions.
@@ A.1. Boundary conditions
The derivations of the transition intensities in Sec. 3.2 assume an infinite grid that approximates $\mathbb{R}^{2}$. In practice, this is not possible and we must instead enforce a finite boundary condition. There are three kinds of boundaries we can enforce with this grid-based set up: 1) absorbing 2) reflecting 3) periodic. In the first case, there is a net flux out of the grid, thus the process does not conserve probability. This creates significant problems when attempting to solve for the stationary distribution. In the second case, a reflective boundary implies ‘no-flux’ at the border. This does preserve probability and is the most sensible choice for unbounded processes on $\mathbb{R}^{2}$. For a cell at the boundary, a transition (or two for the corner cells) is not possible, thus we must correct the diagonal entry of the Laplacian correspondingly in order to conserve probability. In all cases except the Kuramoto oscillator, we opt for this boundary condition. In the case of the Kuramoto oscillator, we have a periodic domain, thus we mirror this with a periodic boundary condition. This preserves probability and transitions are possible from one side of the grid to the other side. The transition intensities must be amended accordingly.
@@ A.2. Solving for the stationary distribution
Solving for the stationary distribution involves solving the equation $L\pi=0$. Whilst for small number of discrete-states this can be calculated numerically in an efficient and exact way, for large state spaces (fine grids) this calculation becomes prohibitively slow. Instead, we solve,
@@
using LSQR [Paige1982LSQR], which capitalises on the sparsity of $L$. Whilst our theoretical derivations confirm that entries of $\pi$ should have the same sign, numerical instabilities may occur. For extremely fine grids, the discretisation error can be comparable to the error incurred through this approximation, thus convergence analysis becomes challenging.
Appendix B Sampling method for diffusions
B.1. Exact sampling for the Ornstein-Uhlenbeck process

### v2

@@ 5.2.1. An application to real-world trajectories: group-polarisation in schooling fish
Finally, we apply this approach to a real-world case of a stationary diffusion, the polarisation trajectories of schooling fish, and determine that such a process is not in a NESS. We follow the analysis of Jhawar et al. [33], who analyse populations of $N=15,30$ and 60 fish moving in a tank. In particular, the positions, $\mathbf{x}_{i}(t)$, and velocities, $\mathbf{v}_{i}(t)$, of the fish can be extracted from video recordings. The group polarisation is an order parameter describing their collective alignment given by
@@
Next, we investigate if the stationary dynamics of the schooling fish are best described as a NESS or an ESS. For each of the population sizes, we infer the EPR of the original trajectory as well as the EPR of 100 surrogate models, obtained using the shuffling procedure. Fig. 18 shows that, across population size, the EPR is not significantly higher for the original trajectories when compared to the surrogate model, from which we conclude that the process is not in a NESS. This result supports a well-known phenomenon in collective dynamics and active matter. Whilst collective dynamics are inherently active at a local level [74], as they dissipate energy in order for agents to adjust their states dynamically, their collective behaviour can be described by equilibrium steady-state dynamics respecting time-reversal symmetry [17] and maximum-entropy distributions [5, 51].
@@ 6. Discussion
Nonequilibrium steady-states are ubiquitous phenomena in physical and biological systems [26, 54, 16]. Approximating the continuous-state dynamics of these diffusions with discrete-state models is a simple and effective way to analyse their dynamics, and has found applications to physical and biological data at multiple scales [2, 40, 42, 53]. Nevertheless, the consistency of the properties of the NESS between the continuous and discrete spaces was not well understood. Moreover, these approaches suffer the complex and obscurant effects of coarse-graining which can hide probability flux leading to underestimates of the entropy production rate [13] as well as induce non-Markovian effects that lead to artefactual scaling relationships [66].
In this paper, we presented a new approach to analytically derive a discrete-state Markov process from stationary planar diffusion, using finite-volume methods. Moreover, we showed that the EPR of the approximation converges in the infinite-state limit, which we illustrated with both solvable and unsolvable examples of stationary diffusions. In the case of unsolvable processes, this approximation gives an accurate way to investigate the behaviour of the EPR under changes to the parameters, which we illustrated with the stochastic van der Pol and Kuramoto models. Finally, we considered the problem of statistical inference from trajectory data. Using sampled paths, our results show that discrete-state models significantly underestimate the EPR of a NESS, even for long trajectories and fine grids. Nevertheless, discrete-state models can be used to define a hypothesis test that can be used to determine if a trajectory is from a NESS or not, which we, again, illustrated with sampled trajectories. Finally, we considered an illustrative application to real-world trajectories by analysing the group-polarisation trajectories from schooling fish [33].
@@
Python code, Jupyter notebooks, and trajectory data are available at https://github.com/rnartallo/coarse_grained_diffusion. The full dataset for the schooling fish has been made available by Jhawar et al. at https://zenodo.org/records/3632470.
@@ Acknowledgements
The authors would like to thank Alexander Strang, Tassilo Schwarz and Jean-Charles Delvenne for important conversations about coarse-graining and stochastic processes. Additionally, they would like to thank Endre Süli, José Antonio Carrillo, Maria Bruna, and Jaime Agudo-Canalejo for their advice on finite-volume discretisations. Finally, they would like to thank Jhawar et al. for making their data on schooling fish publicly available.
@@ Appendix A The Scharfetter-Gummel discretisation
To derive the Scharfetter-Gummel discretisation in Eq. (21), we follow the analysis of Ref. [18], extending to the 2D case with non-constant drift. We begin with the FVA of the FP equation
@@
which is Eq. (19) of the main text. Next, we assume that the flux $J^{x}_{i+\frac{1}{2}}$, drift $f^{x}_{i+\frac{1}{2}}$ and diffusion $D^{x}_{i+\frac{1}{2}}$ are constant over the interval $x_{i},x_{i+1}]$. As a result, $p^{j}(x):=p(x,y_{j})$ satisfies the first-order ODE with constant coefficients
(68) $J^{x}_{i+\frac{1}{2},j}$ $=f^{x}_{i+\frac{1}{2},j}p^{j}(x)-D^{x}_{i+\frac{1}{2},j}\frac{dp^{j}}{dx}(x),$
@@
We then integrate this over the range $x_{i},x_{i+1}]$ where $\Delta x=x_{i+1}-x_{i}$. This yields
@@
Reintroducing the subscripts, we obtain
@@
Whilst not reported here, we found that other structure-preserving schemes such as the Chang-Cooper discretisation [57] yielded valid master equations for these diffusions, whilst the centred-difference scheme only did so under step-size conditions. A number of structure-preserving schemes for the FP equation can be parametrised as a single family as described in Ref. [30].
@@ A.1. Boundary conditions
The derivations of the transition rates in Sec. 3.2 assume an infinite grid that approximates $\mathbb{R}^{2}$. In practice, this is not possible and we must instead enforce a finite boundary condition. There are three kinds of boundaries we can enforce with this grid-based set up: 1) absorbing 2) reflecting 3) periodic. In the first case, there is a net flux out of the grid, thus the process does not conserve probability. This creates significant problems when attempting to solve for the stationary distribution. In the second case, a reflective boundary implies ‘no-flux’ at the border. This does preserve probability and is the most sensible choice for unbounded processes on $\mathbb{R}^{2}$. For a cell at the boundary, a transition (or two for the corner cells) is not possible, thus we must correct the diagonal entry of the Laplacian correspondingly in order to conserve probability. In all cases except the Kuramoto oscillator, we opt for this boundary condition. In the case of the Kuramoto oscillator, we have a periodic domain, thus we mirror this with a periodic boundary condition. This preserves probability and transitions are possible from one side of the grid to the other side. The transition rates must be amended accordingly.
@@ A.2. Solving for the stationary distribution
Solving for the stationary distribution involves solving the equation $L\pi=0$. Whilst for small number of discrete-states this can be calculated numerically in an efficient and exact way, for large state spaces (fine grids) this calculation becomes prohibitively slow. Instead, we solve
@@
using LSQR [55], which capitalises on the sparsity of $L$. Whilst our theoretical derivations confirm that entries of $\pi$ should have the same sign, numerical instabilities may occur. For extremely fine grids, the discretisation error can be comparable to the error incurred through this approximation, thus convergence analysis becomes challenging.
Appendix B Non-diagonal diffusion
In this paper we focus on the case of diagonal diffusion, and use the SG discretisation with a five-point stencil. This implies that the $dp_{i,j}/dt$ depends only on the four adjacent neighbours $\{p_{i+1,j},p_{i-1,j},p_{i,j+1},p_{i,j-1}\}$. Here we will show that using a non-diagonal diffusion matrix requires a nine-point stencil, including the additional terms $\{p_{i+1,j+1},p_{i-1,j+1},p_{i+1,j-1},p_{i-1,j-1}\}$. However, we show that this discretisation does not yield a valid master equation. Finally, we consider the use of shifted grids for coarse-graining processes which are diagonal under a coordinate transform, which includes the case of constant diffusion.
B.1. Discretising non-diagonal diffusion with a nine-point stencil
For simplicity, we will consider an SDE with no drift and non-diagonal, spatially-dependent diffusion, where $D(x,y)$ is given by
(75) $D(x,y)=\begin{pmatrix}D^{x}(x,y)&D^{xy}(x,y)\\ D^{xy}(x,y)&D^{y}(x,y)\end{pmatrix}$,
where $D$ is positive definite. The FP flux is then given by $J=-\nabla\cdot(Dp)$. This process cannot be discretised with a generalisation of the SG discretisation due to these off-diagonal entries producing cross derivatives. We first note that
(76) $\nabla\cdot(Dp)$ $=\begin{pmatrix}\partial_{x}\left(D^{x}p\right)+\partial_{y}\left(D^{xy}p\right)\\ \partial_{x}\left(D^{xy}p\right)+\partial_{y}\left(D^{y}p\right)\end{pmatrix}$.
To obtain an expression for the discrete flux, we will use traditional centred differences [39], for example we have
(77) $\partial_{x}\left(D^{x}p\right)|_{i+\frac{1}{2},j}$ $\approx\frac{D^{x}_{i+1,j}p_{i+1,j}-D^{x}_{i,j}p_{i,j}}{\Delta x},$
(78) $\partial_{y}\left(D^{xy}p\right)|_{i+\frac{1}{2},j}$ $\approx\frac{D^{x}_{i+\frac{1}{2},j+1}p_{i+\frac{1}{2},j+1}-D^{x}_{i+\frac{1}{2},j-1}p_{i+\frac{1}{2},j-1}}{2\Delta x},$
(79) $\approx\frac{D^{x}_{i+\frac{1}{2},j+1}\left(p_{i+1,j+1}+p_{i,j+1}\right)-D^{x}_{i+\frac{1}{2},j-1}\left(p_{i+1,j-1}+p_{i,j-1}\right)}{4\Delta x}.$
Repeating this calculation at each boundary, we can obtain the transition rates between states,
(80) $L_{(i,j)(i+1,j)}$ $=\frac{D^{xx}_{i+1,j}}{\Delta x^{2}}+\frac{D^{xy}_{i+1,j+\frac{1}{2}}-D^{xy}_{i+1,j-\frac{1}{2}}}{4\Delta x\Delta y},$
(81) $L_{(i,j)(i-1,j)}$ $=\frac{D^{xx}_{i-1,j}}{\Delta x^{2}}-\frac{D^{xy}_{i-1,j+\frac{1}{2}}-D^{xy}_{i-1,j-\frac{1}{2}}}{4\Delta x\Delta y},$
(82) $L_{(i,j)(i,j+1)}$ $=\frac{D^{yy}_{i,j+1}}{\Delta y^{2}}+\frac{D^{xy}_{i+\frac{1}{2},j+1}-D^{xy}_{i-\frac{1}{2},j+1}}{4\Delta x\Delta y},$
(83) $L_{(i,j)(i,j-1)}$ $=\frac{D^{yy}_{i,j-1}}{\Delta y^{2}}-\frac{D^{xy}_{i+\frac{1}{2},j-1}-D^{xy}_{i-\frac{1}{2},j-1}}{4\Delta x\Delta y},$
(84) $L_{(i,j)(i+1,j+1)}$ $=\frac{D^{xy}_{i+\frac{1}{2},j+1}+D^{xy}_{i+1,j+\frac{1}{2}}}{4\Delta x\Delta y},$
(85) $L_{(i,j)(i-1,j+1)}$ $=-\frac{D^{xy}_{i-\frac{1}{2},j+1}+D^{xy}_{i-1,j+\frac{1}{2}}}{4\Delta x\Delta y},$
(86) $L_{(i,j)(i+1,j-1)}$ $=-\frac{D^{xy}_{i+\frac{1}{2},j-1}+D^{xy}_{i+1,j-\frac{1}{2}}}{4\Delta x\Delta y},$
(87) $L_{(i,j)(i-1,j-1)}$ $=\frac{D^{xy}_{i-\frac{1}{2},j-1}+D^{xy}_{i-1,j-\frac{1}{2}}}{4\Delta x\Delta y}.$
Examining these rates we see that the positivity constraint is violated e.g. if $L_{(i,j)(i+1,j+1)}>0$ then $L_{(i+1,j)(i,j+1)}<0$. As a result, we cannot define a valid Markov chain with this rectangular grid unless $D^{xy}=0$.
B.2. Diffusion which is diagonal under a coordinate transform
Whilst we have showed that any process with non-diagonal diffusion does not yield a valid master equation, we now consider the special case of diffusion which is diagonal under a coordinate transform. Consider a process where the diffusion matrix can be written as,
(88) $D(x,y)$ $=V\Lambda(x,y)V^{-1},$
where $V\in\mathbb{R}^{d\times d}$ is a matrix whose columns are the eigenvectors of $D$ and $\Lambda(x,y)=\text{diag}(\lambda_{1}(x,y),...,\lambda_{d}(x,y))$ is the diagonal matrix of eigenvalues i.e. the eigenvalues may change with space, but the eigenvectors remain constant. Clearly such a matrix-valued function can be diagonalised, $\Lambda(x,y)=V^{-1}D(x,y)V$, where the original approach can be applied in the transformed coordinates, which we call $(\nu,\eta)$ in $\mathbb{R}^{2}$. This is equivalent to constructing a grid of rotated rectangles where the nodes are placed at
(89) $(x_{i},y_{j})=\mathbf{v}_{1}i\Delta\nu+\mathbf{v}_{2}j\Delta\eta,$
where $\mathbf{v}_{1},\mathbf{v}_{2}$ are the columns of $V$.
Example: Ornstein-Uhlenbeck process
Figure 19. Approximating processes with non-diagonal diffusion. $a)$ We can apply the original approach, using a square grid, in the transformed space where the diffusion is diagonal. We can then return to the original coordinates where the grid is then rotated. $b)$ We illustrate this with the stationary distribution of the approximation and the exact density in both the transformed and original space.
We illustrate this coordinate transform approach on an OU process with
(90) $B$ $=\begin{pmatrix}2&-1\\ 1&2\end{pmatrix},\hskip 10.0ptD=\begin{pmatrix}4&1\\ 1&2\end{pmatrix}$,
i.e. where the diffusion is not diagonal, but diagonalisable and constant. We transform into the coordinate system where the diffusion is diagonal, and apply the SG discretisation with a square grid. The transition rates can then be considered as acting on a rotated grid in the original coordinates. Fig. 19 illustrates the transition rates (Panel $a)$), and stationary distribution (Panel $b)$) compared to the exact density in both the original and transformed coordinates.
For a diffusion matrix which does not have constant eigenvectors, it would be best approximated with an adaptive, irregular grid. However, this severely complicated both theoretical and numerical analysis.
Appendix C A variational approach to nonequilibrium steady-states
Here, we will briefly consider the energy dissipation principle, which is central to variational approaches to NESS [4, 75, 48, 49, 50, 35].
We begin with a diffusive process of the form in Eq. (1), and define the thermodynamic force,
(91) $F(x)$ $=\chi^{-1}(x)J(x)=D^{-1}(x)f_{\text{irr}}(x),$
where $\chi(x)=D(x)p(x)$ is the mobility. This force is the conjugate variable to the flux, allowing us to write the EPR at stationarity in the form $\Phi=\int J_{\text{ss}}\cdot F_{\text{ss}}\;dx$. Crucially, the variational structure is defined by two convex conjugate functionals. First, we define the dissipation potential $\Psi(J)$, which captures the thermodynamic cost of observing a given fluctuation in the flux $J$ [4],
(92) $\Psi(J)$ $=\frac{1}{2}\int J(x)^{\top}\chi^{-1}(x)J(x)\;dx.$
Second, we define the dual dissipation potential, $\Psi^{*}(F)$, which acts on the forces,
(93) $\Psi^{*}(F)$ $=\frac{1}{2}\int F(x)^{\top}\chi(x)F(x)\;dx=\frac{1}{2}\int p(x)F(x)^{\top}D(x)F(x)\;dx.$
The central result of the energy dissipation principle is that the physical steady state minimizes the functional $\Psi(J)+\Psi^{*}(F)-\int J\cdot Fdx$. Consequently, at the physical steady state (where $J_{\text{ss}}=\chi F_{\text{ss}}$), the total entropy production decomposes symmetrically into the sum of the potentials i.e. $\Phi=\Psi(J_{\text{ss}})+\Psi^{*}(F_{\text{ss}})$. It is clear that the EPR has a quadratic relationship with the flux, stemming from the quadratic nature of the dissipation potential.
On the other hand, the dissipation potential of a CTMC is non-quadratic [48, 35]. To show this, we consider a CTMC and write the discrete flux, Eq. (12), as
(94) $J_{ij}$ $=\Lambda_{ij}\sinh\left(\frac{F_{ij}}{2}\right),$
where $\Lambda_{ij}=2\sqrt{L_{ij}\pi_{j}L_{ji}\pi_{i}}$ is the discrete analogue of the mobility, and $F_{ij}=\log\left(L_{ji}\pi_{i}/L_{ij}\pi_{j}\right)$ is the discrete force [35]. At stationarity, we have that $\Phi=\frac{1}{2}\sum_{i,j}J_{ij}F_{ij}$. [^19] The factor of $\frac{1}{2}$ stems from double counting each directed edge $(i,j)$ which has the same contribution to the EPR. To derive the dissipation potential and its conjugate, we note that $J_{ij}=\frac{\partial\Psi^{*}_{ij}}{\partial F_{ij}}$, thus
(95) $\Psi^{*}_{ij}$ $=2\Lambda_{ij}\left(\cosh\left(\frac{F_{ij}}{2}\right)-1\right).$
Via the Legendre transform, we have that
(96) $\Psi_{ij}(J_{ij})=\sup_{F_{ij}}\left(J_{ij}F_{ij}-\Psi^{*}_{ij}(F_{ij})\right),$
which is achieved at $F_{ij}=2\operatorname{arcsinh}\left(\frac{J_{ij}}{\Lambda_{ij}}\right)$, thus we have that [^20] We use the identity $\cosh(\operatorname{arcsinh}(x))=\sqrt{1+x^{2}}$.
(97) $\Psi_{ij}(J_{ij})$ $=2J_{ij}\operatorname{arcsinh}\left(\frac{J_{ij}}{\Lambda_{ij}}\right)-2\Lambda_{ij}\left(\sqrt{1+\frac{J_{ij}^{2}}{\Lambda_{ij}^{2}}}-1\right).$
The EPR along a directed edge $(i,j)$, $\Phi_{ij}$, can be written as the sum
(98) $\Phi_{ij}$ $=\Psi^{*}_{ij}(F_{ij})+\Psi_{ij}(J_{ij}),$
(99) $=2J_{ij}\operatorname{arcsinh}\left(\frac{J_{ij}}{\Lambda_{ij}}\right).$
Clearly, this is function is not quadratic in the flux. However, in the continuum limit that we consider in Sec. 3.3.1, the quantity $J_{ij}/\Lambda_{ij}\rightarrow 0$, as the rates increase. As a result, we can use the expansion $\operatorname{arcsinh}(x)\approx x+O(x^{3})$, to obtain
(100) $\Phi_{ij}$ $\approx\frac{2J_{ij}^{2}}{\Lambda_{ij}},$
where the EPR is a quadratic function of the flux, normalised by the mobility, in direct analogy with the expression for a diffusive process.
Appendix D Sampling method for diffusions
D.1. Exact sampling for the Ornstein-Uhlenbeck process